A radar unit is used to measure the speed of cars at NLEX. NLEX records show that the measures of speed are normally distributed with a mean of 94 km/hr and a standard deviation of 12 km/hr. The speed of the top 15% drivers are considered over speeding and are apprehended. Therefore, for a driver of the car to be charged of over speeding, his speed should be a minimum of Blank 1 km/hr. Round off your answer to two (2) decimal places.



1
Expert's answer
2021-10-06T14:43:26-0400

Let "X=" the speed of car: "X\sim N(\mu, \sigma^2)."

Given "\mu=94, \sigma=12"


"P(X>x)=1-P(X\leq x)"

"=1-P(Z\leq \dfrac{x-\mu}{\sigma})"

"=1-P(Z\leq \dfrac{x-94}{12})=0.15"

"P(Z\leq \dfrac{x-94}{12})=0.85"

"\dfrac{x-94}{12}\approx1.036433"

"x=94+12(1.036433)"

"x=106.44"

His speed should be a minimum of "106.44" km/hr.